Cities A, Band C are equidistant from each other. A motorist travels from A to

B at 30 km/h; from B to C at 40 km/h and from C to A at 50 km/h. Determine

his average speed for the entire trip.


1
Expert's answer
2021-07-11T17:54:22-0400

Let "AB=BC=CA=x\ km."

The total distance will be "AB+BC+CA=3x."

On side "AB" the time will be "t_1=\dfrac{x}{30}\ h."


On side "BC" the time will be "t_2=\dfrac{x}{40}\ h."


On side "CA" the time will be "t_3=\dfrac{x}{50}\ h."

The total time of the trip will be


"t_1+t_2+t_3=\dfrac{x}{30}+\dfrac{x}{40}+\dfrac{x}{50}=\dfrac{47x}{600}"

The average speed for the entire trip is


"v_{ave}=\dfrac{3x}{\dfrac{47x}{600}}=\dfrac{1800}{47} (km/h)\approx38.298\ km/h"